OpenStudy (anonymous):

when x equals 8, y equals 20. find y when x equals 42

OpenStudy (anonymous):

105

OpenStudy (anonymous):

simplest equation that would fit this is y = ax where a is some number we dont know 20 = a x 8 solve for a a = 20/8 = 2.5 plug that value of a back in and use your new points y = 2.5 x 42 = 105

OpenStudy (anonymous):

because 5.25*8=42 so do 2o *5.25

OpenStudy (confluxepic):

@JamesWolf is correct. Great job for showing the work.

OpenStudy (anonymous):

hey i answered first and showed work

OpenStudy (anonymous):

the only diff is that mine took 2 messages because of my dumb computor @confluxepic

OpenStudy (confluxepic):

Don't devalue the question and answer process! Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!